OpenStudy (anonymous):

Solve inequality -n-4<3

OpenStudy (anonymous):

-x-4 = 3 -4-3 = x -x-7 = 0 the equation has to be less than 3, knowing x=-7 makes the equation 0 tells us that everything everything close to -7 to positive infinity is in the domain if 4-n<3. Thus the domain would be (-7, +infinity)